What is the mass of a 15 centimeter cube of iron if the density of iron is 7.87 grams/centimeter³?
1
Expert's answer
2019-12-27T02:12:53-0500

Density, "d= m\/v"

m=15 *7.87 =118.05g

V= 15cm3

d=118.05/15

d=7.87g/cm3